Solve for x if (g ◦ f)(x) = 1. Here, f(x) = (xlog(x) · x2) and g(x) = log(x) + 1.



1
Expert's answer
2021-10-12T10:31:28-0400

Solution:

 f(x) = (xlog(x) · x2) and g(x) = log(x) + 1.

"(gof)(x)=g(f(x))=g(x^{\log x} · x^2)
\\=g(x^{2+\log x})
\\=\log (x^{2+\log x})+1"

Now, "(gof)(x)=1"

"\Rightarrow \log (x^{2+\log x})+1=1
\\\Rightarrow \log (x^{2+\log x})=0
\\\Rightarrow x^{2+\log x}=1
\\\Rightarrow x^{2+\log x}=x^0
\\ \\\Rightarrow {2+\log x}=0
\\ \\\Rightarrow \log x=-2
\\\Rightarrow x=e^{-2}"

Hence, real solution is "x=e^{-2}"